# Multimodal-enhanced Federated Recommendation: A Group-wise Fusion Approach

### v1

@@ 4.3. Preference-Guided Distillation
To effectively incorporate multimodal features, we propose a preference-guided distillation strategy that projects group-level multimodal features into the preference space, providing auxiliary supervision for client-side training to enable goal-aligned knowledge transfer and enhanced personalization.
Map group-level multimodal features into preference space. Given the group-level multimodal features $\{Q_{l}\}_{l=1}^{g}$, we map them into the preference space via aggregated group prediction functions. Specifically, we obtain the group prediction function by averaging client functions within each group, $\overline{P}_{l}=\text{AvgPooling}(\{P_{u}\}_{u\in G_{l}})$, and apply $\overline{P}_{l}$ to $Q_{l}$ to generate the group’s preference representation,
@@
Once the group’s preference representations $\{E_{\mathcal{M}\_l}\}_{l=1}^{g}$ are obtained, the server would distribute them to clients for improving the local model’s training.
Distill group preference to client recommendation models. Each client trains its recommendation model locally on private data. Upon receiving the group-level preference representation from the server, it incorporates this as auxiliary supervision to refine the local model. This distillation transfers multimodal and group-aligned knowledge to enhance semantic capture while preserving personalization. Formally, the distillation objective for client $u$ in group $l$ is,
@@
where $\theta_{u}$ denote the $u$-th client’s model parameters, and $\mathcal{L}_{dis}$ is the distillation loss function, e.g., the KL divergency between model’s prediction $\hat{\mathcal{R}}_{u}$ and group preference $E_{\mathcal{M}\_l}$.
@@ 4.4. Optimization Objective and Algorithm
Optimization objective. GFMFR is a multimodal FR system that distills group-aware preference signals from multimodal item semantics to guide personalized client model training. Based on the above procedure, we reformulate the overall optimization objective in Eq. (2) as follows,
@@
where $\lambda$ is the coefficient of distillation loss. $(E_{G}^{*},\phi^{*})$ are the optimized parameters of the multimodal aggregation module, used to generate group-level multimodal representations $Q_{l}$, which are subsequently mapped to preference representations $E_{\mathcal{M}\_l}$. Here, $l$ denotes the group index assigned to the $u$-th client, which is dynamically updated during the federated optimization process. The detailed GFMFR procedure is illustrated in Appendix.
Algorithm. To solve the above optimization objective, we conduct an iterative algorithm between the server and clients to learn the multimodal FR system. As illustrated in Appendix, before the federated optimization begins, the server first performs a unified initialization of the globally shared parameters, including the item embedding $E_{\mathcal{I}}$ and the prediction function $P$. In parallel, high-quality initial multimodal representations $\{\mathcal{E}_{m}\}_{m=1}^{k}$ are obtained using a powerful multimodal representation learning model. In each communication round $\bm{t}$, selected clients $u\in B^{(t)}$ initialize local models with the latest shared parameters $E_{\mathcal{I}}^{(t)}$ and $P^{(t)}$, perform local updates using personal data $\mathcal{R}_{u}$, and upload the updated shared components to the server. Upon receiving the updated shared parameters $\{E_{\mathcal{I}}^{u}\}_{u\in B^{(t)}}$ and $\{P_{u}\}_{u\in B^{(t)}}$, the server first performs global aggregation to obtain the new shared parameters $E_{\mathcal{I}}^{(t+1)}$ and $P^{(t+1)}$. Concurrently, it clusters clients into groups $\{G_{1},G_{2},...,G_{g}\}^{(t)}$ based on the shared prediction functions $\{P_{u}\}_{u\in B^{(t)}}$, and trains a multimodal aggregation module with group-specific item embeddings $\{C_{l}^{(t)}\}_{l=1}^{g}$ to derive group-level multimodal representations $\{Q_{l}^{(t)}\}_{l=1}^{g}$. These representations are then transformed into group preference signals $\{E_{\mathcal{M}\_l}^{(t)}\}_{l=1}^{g}$ via the group-specific prediction function $\{\overline{P}_{l}^{(t)}\}_{l=1}^{g}$, which are subsequently distributed to corresponding clients $u\in B^{(t)}$ to refine their local recommendation models with the distillation loss $\mathcal{L}_{dis}$.
4.5. Discussion about Practical Applicability
Usage guideline. The proposed multimodal-enhanced FR framework emphasizes modularity and compatibility for seamless integration into existing FR pipelines. The multimodal representation learning is decomposed into three stages for practical deployment and customization. First, initial modality-specific representations can be obtained by directly using powerful pre-trained models on the server or trained from scratch, supporting flexible domain adaptation. Second, the group-aware multimodal aggregation module supports flexible grouping strategies (e.g., behavior, content, geography) and aggregation mechanisms (e.g., MLP (Rosenblatt, 1958), self-attention (Vaswani et al., 2017), mixture-of-experts (Jacobs et al., 1991)), with adjustable aggregation frequency balancing personalization and efficiency. Third, the preference-guided distillation can be replaced by alternative integration methods (e.g., adapter modules, attention fusion) to align multimodal knowledge with local objectives. These design choices ensure adaptability across diverse system constraints and scenarios.
Efficiency analysis. Multimodal item attributes, derived from raw data such as images, text, and audio, require high-capacity models for effective encoding. In GFMFR, storing and encoding these attributes is fully offloaded to the server, leveraging its superior computational and storage resources. This eliminates the burden on resource-limited clients from handling large multimodal data or complex models. The server further compresses rich multimodal semantics into low-dimensional preference representations (e.g., dimension 2 in implicit feedback tasks), which are then transmitted to clients to enhance local training. This approach preserves task-relevant semantics while significantly reducing communication overhead in federated optimization.
Security analysis. Unlike conventional ID-based representations, multimodal features are semantically rich but may inadvertently encode sensitive user attributes when learned locally, increasing the risk of privacy breaches via gradient inversion or attribute inference. In our framework, representation learning is entirely server-side, and clients receive only compact item preference vectors derived from multimodal features, which are less informative and more resilient to such attacks. This design mitigates privacy risks associated with multimodal signals in FR while remaining compatible with ID-based architectures, and amenable to integration with techniques like Differential Privacy (Choi et al., 2018) and Homomorphic Encryption (Acar et al., 2018). Further analysis is provided in the experiments.
5. Experiment
5.1. Experimental Setup
Dataset. We conduct experiments on five multimodal recommendation benchmark datasets from the Amazon collection(Hou et al., 2024), each representing a different item category (Beauty, Tools_and_Home, Toys_and_Games, Digital_Music, Office_Products), to evaluate the effectiveness of our proposed method. Detailed data descriptions are summarized in Appendix.
Baselines. Our proposed GFMFR is a general multimodal FR framework that can be seamlessly integrated with existing ID-based FR models. To verify its versatility, we select three advanced ID-based FR models as the backbones, including FedNCF (Perifanis and Efraimidis, 2022), PFedRec (Zhang et al., 2023a) and FedRAP (Li et al., 2024b). In addition, we include three existing multimodal FR models as baselines for comparison, including FedMMR, AMMFRS and FedMR. In addition, we extend a centralized multimodal recommendation model (Zhang et al., 2017) to the federated setting (named FedMLP) by employing an MLP-based fusion strategy on the client side, thereby enriching the comparison pool with a comprehensive coverage of commonly used multimodal fusion techniques. More details can be found in Appendix.
Evaluation protocol and implementation details. We evaluate the performance using two ranking-based metrics for top-k recommendation, i.e., Hit Rate (HR@k) and Normalized Discounted Cumulative Gain (NDCG@k). Specifically, we predict over the full item set and calculate the metrics based on the rank of the test item within the top-50 recommendation list. All experiments are conducted five times, with their average results consistently reported in the paper. More implementation details can be found in Appendix.
5.2. Overall Comparison
Overall performance. Table 1 reports HR@50 and NDCG@50 comparisons across five datasets. Key insights from the results are as follows:
(1) Integrating multimodal information into FR models can typically enhance system performance. By incorporating semantic cues, multimodal data enriches the system’s understanding of items. This deeper understanding enables more effective knowledge transfer between clients, which in turn boosts system performance. (2) Our method demonstrates superior performance compared to baselines. Existing multimodal FR methods focus on enhancing ID embeddings via specific fusion strategies, focusing primarily on feature enhancement. In contrast, our method leverages multimodal information to learn preference representations that are more aligned with the optimization objectives of recommendation models, leading to improved performance. (3) Our method showcases exceptional compatibility across various FR backbones. GFMFR serves as a general multimodal FR framework that can be seamlessly integrated with existing ID-based methods. Experimental results demonstrate that incorporating our model into three different FR backbones consistently yields the best performance across all datasets, highlighting its strong compatibility.
@@ 5.2. Overall Comparison
Efficiency comparison. We further evaluate and compare the efficiency of our method and baselines from three aspects: (1) client-side storage overhead, including multimodal data representations (preferences) and recommendation model parameters; (2) communication overhead, covering the initial distribution of multimodal representations and the per-round transmission cost; and (3) per-epoch training time of the client-side recommendation model. As shown in Appendix, baselines AMMFRS, FedMR and FedMLP incur higher client storage overhead due to the need to store raw multimodal features and the integration of fusion modules within local models. Baseline FedMMR’s storage overhead primarily stems from its architecture by integrating an elaborate contrastive loss. These baselines also require longer client training time due to more complex local models. In terms of communication, baseline AMMFRS, FedMR and FedMLP transmit full initial multimodal representations to all clients before training, resulting in substantial communication overhead. In contrast, our method only exchanges low-dimensional group preference vectors (e.g., dimension 2 vs. 512), introducing minimal additional overhead. Overall, our method achieves higher efficiency and is better suited for real-world deployment constraints.
@@ 5.3. Ablation Study
This subsection examines the impact of the group-aware multimodal aggregation mechanism and the preference-guided distillation strategy on model performance. The analysis clarifies their individual contributions and demonstrates how each component enhances overall effectiveness. We conduct the evaluation by incorporating our method into PFedRec and reporting HR@50 results across all datasets. For results on NDCG@50, see Figure LABEL:fig:ablation_ndcg in Appendix.
Group-aware multimodal aggregation mechanism. We perform a step-wise ablation study with three key variants: replace preference-based clustering with random grouping to assess grouping quality (RandomGroup); treat all users as one group to examine the need for group differentiation (SingleGroup); and use group-specific aggregation modules instead of a shared one to evaluate the benefit of shared representation learning (MultipleAgg). As shown in Figure 2 (a), random grouping notably degrades performance, highlighting the importance of behavior-aware user partitioning. Treating all users as a single group also impairs results, as it overlooks individual preferences. Furthermore, using multiple aggregation modules reduces performance due to increased complexity and the absence of shared representation learning, compromising generalization and stability.
Preference-guided distillation strategy. The preference-enhanced distillation strategy leverages group-level preference representations to improve client training. To examine the effect of preference signal utilization, we compare four coefficient scheduling methods for the distillation loss: (1) match the scale of recommendation loss (ScaleAlign), (2) smooth with the previous round’s coefficient (SmoothCoef), (3) progressively increase the coefficient (ProgressiveScale), and (4) our strategy combines (2) and (3) (Ours). As shown in Figure 2 (b), our method achieves the best performance by balancing training stability and progressive preference integration.
Figure 2. Ablation study results of metric HR@50 on all datasets.
@@ 5.4. Hyper-Parameter Analysis
This subsection investigates the impact of several key hyper-parameters on model performance, including the number of user groups $g$, the number of local training epochs per client $H$, and the client sampling ratio per round $\beta$. For the number of user groups $\bm{g}$, experimental results reveal that model performance is sensitive to the number of user groups, with the best performance varying across datasets. An appropriate group number enables effective user partitioning and better captures preference diversity. For the number of local training epochs $\bm{H}$, setting local epochs to 5 generally yields the best performance. Fewer epochs may lead to underfitting, while excessive training can cause overfitting due to data heterogeneity. For the client sampling ratio $\bm{\beta}$, model performance remains generally comparable across different sampling ratios. Lower ratios may lead to slower convergence, whereas higher ratios help accelerate training and improve resource efficiency. Complete results and analysis on all datasets can be found in Appendix.

### v2

@@ 4.3. Preference-Guided Distillation
To effectively leverage multimodal signals, we propose a preference-guided distillation strategy. This approach projects group-level multimodal features into the user preference space, providing auxiliary supervision during client-side training. In this way, clients incorporate multimodal signals directly aligned with recommendation objectives, enhancing the learning of preference-relevant representations and enabling more personalized recommendations.
4.3.1. Map group-level multimodal features into preference space.
After obtaining group-level multimodal features $\{Q_{l}\}_{l=1}^{g}$, the next step is to project them into the preference space to make them usable for client-side training. This projection allows the multimodal representations, which encode rich item information, to be aligned with how users express preferences in each group. Specifically, we obtain the group prediction function by averaging client functions within each group, $\overline{P}_{l}=\text{AvgPooling}(\{P_{u}\}_{u\in G_{l}})$. This function captures the overall decision patterns of users in the group, reflecting their shared tendencies in ranking and evaluating items. We then apply $\overline{P}_{l}$ to the group-level multimodal features $Q_{l}$ to generate the group’s preference-aligned representation,
@@
Once the group’s preference representations $\{E_{\mathcal{M}\_l}\}_{l=1}^{g}$ are obtained, the server will distribute them to the clients to improve local model’s training.
4.3.2. Distill group preference to client recommendation models.
Each client trains its recommendation model locally on private data. Upon receiving the group-level preference representation from the server, it incorporates this as auxiliary supervision to refine the local model. This distillation guides the client model to align its predictions with group-level user preferences, thereby directly shaping the learning of recommendation-relevant behaviors rather than merely fusing features. Formally, the distillation objective for client $u$ in group $l$ is,
@@
where $\theta_{u}$ denote the $u$-th client’s model parameters, and $\mathcal{L}_{dis}$ is the distillation loss function, e.g., the KL divergency between the client’s predicted outputs $\hat{\mathcal{R}}_{u}$ and the group-level preference representation $E_{\mathcal{M}\_l}$. By leveraging these group-level representations, client models incorporate multimodal signals consistently with group tendencies, facilitating knowledge transfer across users with similar preferences while maintaining computational efficiency.
Algorithm 1 Group-wise Fusion for Multimodal-enhanced Federated Recommendation
ServerProcedure:
1: Initialize globally shared item embedding $E_{\mathcal{I}}^{(1)}$ and prediction function $P^{(1)}$
2: Obtain initial multimodal representations $\{\mathcal{E}_{m}\}_{m=1}^{k}$ with a powerful model
3:for each round $t=1,2,\dots,T$ do
4:  Randomly select a client subset $B^{(t)}$ from $N$ clients based on sampling ratio $\beta$
5:  for client $u\in B^{(t)}$ in parallel do
[…]
7:  end for
8:  $E_{\mathcal{I}}^{(t+1)},P^{(t+1)}\leftarrow\text{GlobalAgg}(\{E_{\mathcal{I}}^{u}\}_{u\in B^{(t)}},\{P_{u}\}_{u\in B^{(t)}})$;
9:  Cluster clients into groups $\{G_{1},G_{2},...,G_{g}\}^{(t)}$ based on $\{P_{u}\}_{u\in B^{(t)}}$ with Eq. (2)
10:  Train a multimodal aggregation module with group-specific item embedding with Eq. (4)
11:  Obtain the group-level multimodal representations $\{Q_{l}^{(t)}\}_{l=1}^{g}$ with Eq. (3)
12:  Calculate the group preference representations $\{E_{\mathcal{M}\_l}^{(t)}\}_{l=1}^{g}$ via the group-specific prediction function with Eq. (5)
13:  Distribute $\{E_{\mathcal{M}\_l}^{(t)}\}_{l=1}^{g}$ to client $i\in B^{(t)}$ to refine local model with Eq. (6). (with $\lambda$ balancing the contribution of distillation loss against the local recommendation loss)
14:end for
[…]
1:if $t=1$ then
2:  Initialize user embedding $E_{\mathcal{U}}$
3:else
4:  Load user embedding from the last round
5:end if
6: Initialize $(E_{\mathcal{I}},P)$ with $(E_{\mathcal{I}}^{(t)},P^{(t)})$
7:for $h=1,2,\dots,H$ do
8:  Update $E_{\mathcal{U}},E_{\mathcal{I}},P$ based on personal data $\mathcal{R}_{u}$ with recommendation loss $\mathcal{L}_{rec}$
9:end for
10:Return $E_{\mathcal{I}},P$
@@ 4.4. Optimization Objective and Algorithm
4.4.1. Optimization objective.
The goal of GFMFR is to jointly learn group-level multimodal representations and guide client models with group-aligned preference signals in the FR framework. We formulate the overall optimization objective as follows,
@@
where $\lambda$ is the coefficient of distillation loss. $(E_{G}^{*},\phi^{*})$ are the optimized parameters of the multimodal aggregation module, used to generate group-level multimodal representations $Q_{l}$, which are subsequently mapped to preference representations $E_{\mathcal{M}\_l}$. Here, $l$ denotes the group index assigned to the $u$-th client, which is dynamically updated during the federated optimization process.
@@
4.4.2. Algorithm.
To solve the above optimization objective, we conduct an iterative algorithm between the server and clients to learn the multimodal FR system. As illustrated in Algorithm 1, before the federated optimization begins, the server first performs a unified initialization of the globally shared parameters, including the item embedding $E_{\mathcal{I}}$ and the prediction function $P$. In parallel, high-quality initial multimodal representations $\{\mathcal{E}_{m}\}_{m=1}^{k}$ are obtained using a powerful multimodal representation learning model. In each communication round $\bm{t}$, selected clients $u\in B^{(t)}$ initialize local models with the latest shared parameters $E_{\mathcal{I}}^{(t)}$ and $P^{(t)}$, perform local updates using personal data $\mathcal{R}_{u}$, and upload the updated shared components to the server. Upon receiving the updated shared parameters $\{E_{\mathcal{I}}^{u}\}_{u\in B^{(t)}}$ and $\{P_{u}\}_{u\in B^{(t)}}$, the server performs global aggregation to obtain the new shared parameters $E_{\mathcal{I}}^{(t+1)}$ and $P^{(t+1)}$. Concurrently, it clusters clients into groups $\{G_{1},G_{2},...,G_{g}\}^{(t)}$ based on the shared prediction functions $\{P_{u}\}_{u\in B^{(t)}}$, and trains a multimodal aggregation module with group-specific item embeddings $\{C_{l}^{(t)}\}_{l=1}^{g}$ to derive group-level multimodal representations $\{Q_{l}^{(t)}\}_{l=1}^{g}$. These representations are then transformed into group preference signals $\{E_{\mathcal{M}\_l}^{(t)}\}_{l=1}^{g}$ via the group-specific prediction function $\{\overline{P}_{l}^{(t)}\}_{l=1}^{g}$, which are subsequently distributed to corresponding clients $u\in B^{(t)}$ to refine their local models with the distillation loss $\mathcal{L}_{dis}$.
4.5. Discussion about Practical Applicability
4.5.1. Usage guideline.
The proposed multimodal-enhanced FR framework emphasizes modularity and compatibility for seamless integration into existing FR pipelines. The multimodal representation learning is decomposed into three stages for practical deployment and customization. First, initial modality-specific representations can be obtained by using powerful pre-trained models on the server or trained from scratch, supporting flexible domain adaptation. Second, the group-aware multimodal aggregation module supports flexible grouping strategies (e.g., behavior, geography) and aggregation mechanisms (e.g., MLP (Rosenblatt, 1958), self-attention (Vaswani et al., 2017), mixture-of-experts (Jacobs et al., 1991)), with adjustable aggregation frequency balancing personalization and efficiency. Third, the preference-guided distillation can be replaced by alternative integration methods (e.g., adapter modules, attention fusion) to align multimodal knowledge with local objectives. These design choices ensure adaptability across diverse system constraints and scenarios.
4.5.2. Efficiency analysis.
Multimodal item attributes, derived from raw data such as images, text, and audio, require high-capacity models for effective encoding. In GFMFR, storing and encoding these attributes is fully offloaded to the server, leveraging its superior computational and storage resources. This eliminates the burden on resource-limited clients from handling large multimodal data or complex models. The server further compresses rich multimodal semantics into low-dimensional preference representations (e.g., dimension 2 in implicit feedback tasks), which are then transmitted to clients to enhance local training. This approach preserves task-relevant semantics while significantly reducing communication overhead in federated optimization.
4.5.3. Security analysis.
Unlike conventional ID-based representations, multimodal features are semantically rich but may inadvertently encode sensitive user attributes when learned locally, increasing the risk of privacy breaches via gradient inversion or attribute inference. In our framework, representation learning is entirely server-side, and clients receive only compact item preference vectors derived from multimodal features, which are less informative and more resilient to such attacks. This design mitigates privacy risks associated with multimodal signals in FR while remaining compatible with ID-based architectures, and amenable to integration with techniques like Differential Privacy (Choi et al., 2018) and Homomorphic Encryption (Acar et al., 2018). See experiments for further analysis.
5. Experiment
5.1. Experimental Setup
Dataset.
We evaluate our method on five multimodal recommendation datasets from Amazon (Hou et al., 2024), including Beauty, Tools_and_Home (TH), Toys_and_Games(TG), Digital_Music(DM), Office_Products(OP). Detailed data descriptions are summarized in Appendix A.1.
Baselines.
Our proposed GFMFR is a general multimodal FR framework that can be seamlessly integrated with existing ID-based FR models. To verify its versatility, we select three advanced ID-based FR models as the backbones, including FedNCF (Perifanis and Efraimidis, 2022), PFedRec (Zhang et al., 2023a) and FedRAP (Li et al., 2024b). In addition, we include three existing multimodal FR models as baselines for comparison, including FedMMR(Li et al., 2024a), AMMFRS(Feng et al., 2024) and FedMR(Li et al., 2024c). We also extend a centralized multimodal recommendation model (Zhang et al., 2017) to the federated setting (named FedMLP) by employing an MLP-based fusion strategy on the client side, thereby enriching the comparison pool with a comprehensive coverage of commonly used multimodal fusion techniques. More details can be found in Appendix A.2.
Evaluation protocol and implementation details.
We evaluate the performance using two ranking-based metrics for top-k recommendation, i.e., Hit Rate (HR@k) and Normalized Discounted Cumulative Gain (NDCG@k). Details can be found in Appendix A.3.
Figure 2. Ablation Study of Group-aware Multimodal Aggregation Mechanism.
Figure 3. Ablation Study of Preference-guided Distillation Strategy.
5.2. Overall Comparison
Overall performance.
Table 1 reports HR@50 and NDCG@50 comparisons across five datasets. Key insights from the results are as follows: (1) Integrating multimodal information into FR models can typically enhance system performance. Effective integration of semantic cues in multimodal data is crucial for the system to achieve a deeper understanding of items. This deeper understanding allows for a more accurate characterization of items, which in turn enhances system performance. (2) Our method demonstrates superior performance compared to baselines. Existing multimodal FR methods primarily generate a unified item representation by leveraging specific fusion strategies to enhance features. In contrast, our method leverages multimodal information to learn preference representations that are more aligned with the optimization objectives of recommendation models, leading to improved performance. (3) Our method showcases exceptional compatibility across various FR backbones. GFMFR serves as a general multimodal FR framework that can be seamlessly integrated with existing ID-based methods. Experimental results demonstrate that incorporating our model into three different FR backbones consistently yields the best performance across all datasets, highlighting its strong compatibility.
Efficiency comparison.
We further evaluate and compare the efficiency of our method and baselines from three aspects: (1) client-side storage overhead, including multimodal data representations (preferences) and recommendation model parameters; (2) communication overhead, covering the initial distribution of multimodal representations and the per-round transmission cost; and (3) per-epoch training time of the client-side recommendation model. Our method consistently achieves lower storage and communication costs as well as faster per-epoch training, demonstrating superior overall efficiency. More details can be found in Appendix B.1.
@@ 5.3. Ablation Study
This subsection examines the impact of the group-aware multimodal aggregation mechanism and the preference-guided distillation strategy on model performance. The analysis clarifies their individual contributions and demonstrates how each component enhances overall effectiveness. We evaluate our method by incorporating it into PFedRec and report the HR@50 and NDCG@50 results across all datasets.
Group-aware multimodal aggregation mechanism.
We perform a step-wise ablation study with three key variants: replace preference-based clustering with random grouping to assess grouping quality (RandomGroup); treat all users as one group to examine the need for group differentiation (SingleGroup); and use group-specific aggregation modules instead of a shared one to evaluate the benefit of shared representation learning (MultipleAgg). As shown in Figure 2, random grouping notably degrades performance, highlighting the importance of behavior-aware user partitioning. Treating all users as a single group also impairs results, as it overlooks individual preferences. Furthermore, using multiple aggregation modules reduces performance due to increased complexity and the absence of shared representation learning, compromising generalization.
Preference-guided distillation strategy.
The preference-guided distillation strategy leverages group-level preference representations to refine client training. To examine the effect of preference signal utilization, we compare four scheduling strategies for the coefficient $\lambda$ (as defined in Eq. 7) of the distillation loss: (1) match the scale of recommendation loss (ScaleAlign), (2) smooth with the previous round’s coefficient (SmoothCoef), (3) progressively increase the coefficient (ProgressiveScale), and (4) our strategy combines (2) and (3) (Ours). As shown in Figure 3, our method achieves the best performance by effectively balancing training stability and the progressive integration of user preferences.
@@ 5.4. Hyper-Parameter Analysis
This subsection investigates the impact of several key hyperparameters on model performance, including the number of user groups $g$, the number of local training epochs per client $H$, and the client sampling ratio per round $\beta$. For the number of user groups $\bm{g}$, experimental results reveal that model performance is sensitive to the number of user groups, with the best performance varying across datasets. An appropriate group number enables effective user partitioning and better captures preference diversity. For the number of local training epochs $\bm{H}$, setting local epochs to 5 generally yields the best performance. Too few local epochs result in underfitted and noisy updates, whereas excessive training can induce overfitting and escalate computational costs. For the client sampling ratio $\bm{\beta}$, model performance remains generally comparable across different sampling ratios. Lower ratios may lead to slower convergence, whereas higher ratios help accelerate training and improve resource efficiency. Complete results and analysis on all datasets can be found in Appendix B.2.